OpenStudy (anonymous):

Add and answer in simplest fractional form. -4/3 + 7/8

OpenStudy (anonymous):

do sumthing by urself

OpenStudy (anonymous):

is it 3/5

OpenStudy (anonymous):

no try once more

OpenStudy (anonymous):

no jimmy dont give the answer

OpenStudy (anonymous):

-11/24

OpenStudy (anonymous):

- 4/3 = 32/24 + 7/8 = 21/24 - this should help

OpenStudy (anonymous):

great u did it !!!

OpenStudy (anonymous):

you're right

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

that's the sprit keep it up and try to solve by urself

OpenStudy (anonymous):

yes - u did it before i posted my reply - by the way i made a mistake - should be -32/24